A Complete Bibliography of The Bell System Technical Journal, 1920–1929

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 February 2022 Version 1.05

Title word cross-reference

Absorption [WB28]. Abstracts [Ano24a, Ano24b, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano26c, Ano26d, Ano26b, Ano27a, Ano27b, Ano27c, Ano27d, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano29b, Ano29c, Ano29d]. Acoustic [Bos29, Mas27, WB28]. Adapt [Kir23b]. Address [GM23, MC23]. Adopted [Smi24]. Adsorption [Ger29]. Advances [Dar23, Dar24c, Dar24a, Dar24d, Dar24b, Dar25a, Dar25b, Dar25c, Dar25d, Dar26a, Dar26b, Dar27a, Dar27b, Dar27c, Dar28a, Dar28b, Dar29a, Dar29b, Fre29]. Affecting [Nyq24]. Airways [Cra28]. Alkali [Ive26]. Alloys [Elm29a]. Along [Car26b, Car28b].

Alternating [Mea25]. Aluminum [Sie29]. American [Ano28i, GK26]. Amplifiers [Cur27, FJ24, PE27, WM26]. Analysis [Bat23, CM22, Hor24, She24, Weg22]. Analyzer [Lan27, MC27, WM24]. Antenna [Fos26]. Apparatus [KH28, Kir23c, Luc24, SF28, Tow29]. Appearing [Ano24a, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano27a, Ano27b, Ano27c, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a]. Application [Cam28, Cur27, Esp22, Ger29, Luc24, Mol27, Mol29, Ros23, SF26, Smi24]. Applications [Har24, Kin23, Mas27, She24, Tho26]. Applied [Mol22, Osb29]. Approximate [Car28b]. Areas [CMC23, Esp27]. Arrays [Fos26]. Articles [Ano29b, Ano29c, Ano29d].

Articulation [FS29]. Asymptotic [Wis29]. Atom [Dar25c, Dar25d, Dar26b]. Atom-Model [Dar25c, Dar25d, Dar26b]. Audition [Fle23, Weg22]. Automatic [Clo27, Rob28]. Average [MC24]. Averages [She26a].

Bands [Har23]. Based [MH26]. Bearing [Fle23]. Before [Dar26b]. Bell [Ano24a, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano26c, Ano26d, Ano26b, Ano27a, Ano27b, Ano27c, Ano27d, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano29b, Ano29c, Ano29d, Cha25, Kir23a, Nan24, Ros23, SP24]. Between [Hel22]. Binaural [HF22]. Binomial [Mol29]. Books [Ano26b]. Braun [Joh29]. Bridge [Fer27, Her28, Sha27, Wil26]. Brief [Dar26b]. Broadcast [BMP26, Esp27]. Building [Car24a]. Building-up [Car24a].

Cable [Buc25, Buc28, CG22, Car26b, Cla23, Dem23, Har28a, Pil22]. Cables [Clo27, Cur27, EH27, Gil26, Gil27, Hit26]. Calculus [Car22, Car25a, Car26b, Car26a]. Capacitance [FB28]. Capacity [Cam22a, WT28, WM26]. Carrier [ADG28, Har23, Hit26, Ros23, Sha27, Wol25]. Carrier-Current [Hit26]. Carrying [WM26]. Cathode [Joh22]. Cell [Ive26]. Central [GW29]. Certain [Nyq24]. Change [Dar29a]. Chapter [Car26b, Car26a]. Characteristics [CG22, Cur25, Gil27, Mil24, Weg22, Zob24]. Charts [She26c]. Chemical [HS23]. Chestnut [Smi25]. Chicago [Pil22]. Circuit [Car25a, Car26b, Car26a, Car27, WS27]. Circuits [Cam23a, Car25b, Cla23, Cri25a, Dem23, Fos24b, Lle26, Nan24, Osb29, SF26, Tru22, Wil26, Zob28]. City [Esp27]. Classical [Dar28a, Dar28b]. Clock [Fer24]. Clock-Controlled [Fer24]. Cobalt [Elm29a]. Coefficient [WB28].

Coefficients [Lac29]. Combination [Mas27]. Communication [Cra28, GK26, Hit26, Osb29, SF28, Tru22]. Compensating [Hoy24]. Complex [HF22, Lan27]. Composite [Zob23]. Composition [HS23]. Condenser [Sie29]. Conditions [MC24]. Conductors [Eng28, Mea25]. Confirmation [Fle26]. Considerations [Bos29]. Constant [Fer24, Hoc26, Zob28]. Constants [Fle25]. Construction [Kir23b, Kir23c]. Containing [Car25c]. Contemporary [Dar23, Dar24c, Dar24a, Dar24d, Dar24b, Dar25a, Dar25b, Dar25c, Dar25d, Dar26a, Dar26b, Dar27a, Dar27b, Dar27c, Dar28a, Dar28b, Dar29a, Dar29b]. Contributions [Ano29e, Ano29f]. Contributors [Ano22a, Ano22b, Ano23a, Ano23b, Ano23c, Ano23d, Ano24c, Ano24d, Ano24e, Ano24f, Ano25e, Ano25f, Ano25g, Ano25h, Ano26e, Ano26f, Ano26g, Ano26h, Ano27e, Ano27f, Ano27g, Ano27h, Ano28e, Ano28f, Ano28g, Ano28h, Ano29g, Ano29h]. Control [She26c, She27, Sto29]. Controlled [Fer24]. Copper [SM27a]. Correction [Mea28, She26b, She26a, Zob28]. Cost [Rho25]. Coverage [Esp27]. Covered [Har28a]. Creosoting [Smi25]. Crossfire [Sha26]. Crystal [Dav28, Lac29]. Crystals [Dar29b]. Current [Car26b, Hit26, Mea25, PK28, Wil26]. Currents [Car24a, CH27]. Curves [Cam23b].

Data [She24, She26b, She26a]. Decibel [Mar29c]. Definite [Mol29]. Densities [Pet28]. Department [And26]. Design [Hoy23, Hoy24, JS25, Mil24, Zob23]. Desired [Tow29]. Determination [Gil27]. Development [SF26]. Developments [Har28a, SM27a]. Deviation [MC24]. Diagrams [Fos26]. Dielectric [Hoc26]. Diffraction [Dar29b, Dav28, Ger29]. Dipole [Wis29]. Direct [Cam22a]. Directive [Fos26]. Distance [ADG28]. Distortion [Her29, Mea28, Zob28]. Distribution [CM22, Hel22, MW29]. Driving [Fos24b]. Driving-Point [Fos24b]. Duplex [Wil26]. Dynamical [CS24, Cra27, Weg22].

Ear [Weg22]. Earth [Car29, NS25]. Economically [Kir23b]. Editor [Ano27i]. Effect [Her29]. Effects [Hoc26, Tru22]. Efficiency [WT28]. Electric [Cam22b, CZ23, Car25a, Car26b, Car26a, Car27, Lan27, NS25, Osb29, Zob23, Zob24]. Electrical [Ano28i, Car28b, FD24, GK26, Gil27, Har24, Mas29, SF28, WM24, Zob28]. Electricity [Dar24b, Dar25a, Dar29c]. Electrode [Hoc26]. Electrolytic [Sie29, WM29]. Electromagnetic [Car27]. Electromagnets [Mil24]. Electron [DG29, Ger29]. Electrons [Dav28, Dav29]. Elementary [Cog28]. Elements [Mas27]. Energy [CM22, Sac25]. Engineering [Cha25, Cog28, Esp22, Pen25, Rho25, She24]. Engineers [Ano28i]. Equipment [Clo27, Dem23, Kir23b]. Errors [She26b, She26a]. Evaluation [Mol29]. Evolved [HS23]. Expansion [Pen29]. Experimental [Fle26]. Exponential [Cam23b]. Extension [NE23]. External [Weg22]. Extraneous [Gil26].

Factor [Hoc26]. Factors [Nyq24]. Fatigue [TG29]. Features [Wat28]. Ferromagnetic [Pet28]. Ferromagnetism [Dar27b]. Ferrous [TG29, TS29]. Fibers [Mas27]. Filter [Cam22b, JS25]. Filters [JS25, Mas27, Zob24, CZ23, Zob23]. Finite [Car26a]. First [Dar25c, Dar28a]. Flux [Pet28]. Fork [Fer24]. Formulas [Wis29]. Forward [Ano22c]. Foundations [Car27]. Fourier [Cam28]. Frequencies [Pet28, Sha27]. Frequency [Dar29a, Fer24, FB28, Fer29, FJ24, Mar29a, MW29, MC27, WM24]. Fundamental [SB26].

Gaging [Rob28]. Gain [Cri25b]. Gas [Ger29]. Gases [Dar25a, HS23]. General [Cha25, Cog28]. Generalization [Car24b, Pen29]. Glasses [HS23]. Gone [Dar26b]. Graphic [Car25c]. Graphical [Bat23, Hor24]. Grid [PK28]. Ground [Car26c, Car29]. Grounded [Cam23a]. Grounding [Tru22].

Harmonic [Pet28]. Hearing [Fle23, Fle25]. Heaviside [Gil25, Car22, Pen29]. Henry [GK26]. High [AE23a, Buc28, Fer29, FJ24, Mar29a, MH26, WT28, Wil22, Wol25]. High-Frequency [Fer29]. High-Speed [Buc28]. Horn [WT28]. Horn-Type [WT28]. Horns [Mas27]. Howling [Fle26]. Humidity [Whe24]. Hysteresigraph [Joh29].

II [Cra27, Dar24c]. III [Dar24a].

Impedance [Car29, Car25c, Cri25b, Cur25, Fos24b, Hoy23, Hoy24, Sha27, WB28]. Impedances [Cam23a, Car26a]. Incomes [Hel22]. Inductance [Fer27, JS25]. Inductive [Car26b, Sha27, Tru22]. Industrial [Cam24]. Influence [WM29]. Information [Har28b]. Inputs [Cur25]. Inspection [DR29]. Institute [Ano28i]. Institution [Ano28i]. Insulating [Hoc22, Hoc26]. Insulation [GW29]. Insulators [WM29]. Integral [Cam28]. Integrals [Mol29]. Interference [Car25b, Gil26]. Interpretation [Fle22]. Intervals [Her28]. Introduction [Dar27c, JS25]. Involved [Bos29]. Ionization [Dar26a]. Iron [Elm29a]. Irregularities [Cri25a, Cri25b]. Issue [Ano22a, Ano22b, Ano23a, Ano23c, Ano23d, Ano24d, Ano24e, Ano24f, Ano25e, Ano25f, Ano25g, Ano25h, Ano26e, Ano26f, Ano26g, Ano26h, Ano27e, Ano27f, Ano27g, Ano27h, Ano28e, Ano28f, Ano28g, Ano28h, Ano29g, Ano29e, Ano29f, Ano29h, Ano23b, Ano24c]. IV [Dar24d]. IX [Car26a, Dar25d].

Joint [Ano28i]. Joseph [GK26]. Journal [Ano25b, Ano25c, Ano25d, Ano26a, Ano27a, Ano27b, Ano27c, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano24a, Ano25a].

Known [HS23].

Labor [Kir23c]. Labor-saving [Kir23c]. Laplacian [Mol29]. Large [CMC23, WT28]. Lead [Har28a]. Lead-Covered [Har28a]. Letters [Ano27i]. Light [Dar28a, Dar28b, Dar29a, Mac29]. Limitation [Cri25b]. Line [Car26a, Wil26]. Linear [Eng28]. Lines [ADG28, Car24a, Hoy23, Hoy24, IHPC25, MC23, Wol25]. Load [WM26]. Loaded [Buc25, Car24a, Clo27, Cri25a, Gil27, Hoy24]. Loading [SF26]. Location [EH27, HF22]. London [WS27]. Long [ADG28, BDW29, Car24a, Cla23, Clo27, Dem23, Nan24]. Long-Wave [BDW29]. Losses [Hoc22]. Loud [Bos29, WT28]. Low [Cur25, Joh22, Pet28].

Machine [CMC23, Rob28]. Magnetic [AE23a, Elm29a, Elm29b]. Maintenance [Har24, Har25, Kir23b]. Making [KH28]. Manufacture [KH28, Pen25, SM27a]. Manufactured [Dod28]. Manufacturing [Har28a]. Map [Kir23a]. Master [MG29]. Material [AE23a, WM29]. Materials [Hoc22, Hoc26, Pet28, WB28]. Mathematics [Cam24]. Matter [Dar29c]. Mean [MW29]. Measurement [Cam22a, Fer27, FB28, Hoc26, SF28, She26b, Siv29, WB28]. Measurements [Bos29, Fer29, Fle23, HS23, Sha27]. Measuring [Her28]. Mechanics [Dar27c]. Meeting [Ano28i]. Men [MC24]. Mesh [Fos24b]. Metal [Ive26]. Metals [TG29, TS29]. Method [Bat23, Dod28, DR29, Hor24, Mas29, Mol29]. Methods [FS29, MH26, She24, TS29]. Metropolitan [CMC23]. Microscopy [Luc24]. Model [Dar25c, Dar25d, Dar26b]. Modes [Lac29]. Modulation [PE27, PK28]. Moisture [WM29]. Morse [Her29]. Most [Kir23b]. Motion [Scr29]. Music [MH26]. Mutual [Cam23a, JS25].

Name [Mar29c]. Natural [Eng28]. Nature [Fle22]. Neighboring [Tru22]. Networks [Car25c, Hoy23, Hoy24, Mas29, Tru22, Zob28]. Neutralization [Sha26]. Nickel [Dav28, Elm29a]. No [Dav27]. Non [Car26b, TG29, TS29]. Non-Ferrous [TG29, TS29]. Non-Inductive [Car26b]. Numerical [Fle25].

Observations [Lac29]. Obtained [She26a]. Obtaining [Mas29]. Ocean [Buc28]. Office [GW29]. Offices [You27]. Oliver [Gil25]. Open [Smi25]. Opens [EH27]. Operating [Wat28]. Operation [Fle26, Lle26]. Operational [Car22, Car25a, Car26b, Car26a]. Order [Kir23b]. Oscillations [CZ23]. Oscillators [Hor24]. Oscillograph [Joh22]. Oscillographs [Mar29b]. Outstanding [Car28a]. Overhead [Car26c]. Owen [Fer27].

Papers [Ano24a, Ano24b, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano26c, Ano26d, Ano26b, Ano27a, Ano27b, Ano27c, Ano27d, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a]. Parallel [CH27, Mea25]. Part [Dar25c, Dar25d, Dar26b, Dar28a, Dar28b, Cra27]. Period [Eng28]. Periodic [CH27]. Periodically [Car24a]. Permalloy [AE23a]. Permeability [AE23a]. Perminvar [Elm29b]. Petersen [Tru22]. Phase [Mea28]. Phenomena [Mar29b]. Philadelphia [Pil22]. Photoelectric [Ive26]. Photomicrography [Luc24]. Physical [Cam22b, Fle23, She24, TS29, Weg22]. Physics [Dar23, Dar24c, Dar24a, Dar24d, Dar24b, Dar25a, Dar25b, Dar25c, Dar25d, Dar26a, Dar26b, Dar27a, Dar27b, Dar27c, Dar28a, Dar28b, Dar29a, Dar29b]. Physics-II [Dar24c]. Physics-IV [Dar24d]. Physics-X [Dar26b]. Physics-XIII [Dar27b]. Picture [Scr29]. Pictures [IHPC25, Sto29]. Pioneer [GK26]. Pittsburgh [Pil22]. Planning [Pen25]. Plant [Kir23c]. Plants [Smi25, You27]. Plates [Lac29]. Point [Fos24b]. Poisson [Cam23b, Tho26]. Polar [Wil26]. Polarization [DG29]. Poles [Smi25]. Police [And26]. Porous [WB28]. Power [Cur25, Hoc22, Hoc26, Sac25, SB26, Siv29, Tru22, WT28, Wil22, Wol25, You27]. Practical [Cam28, Ros23, Smi24]. Practises [Har25]. Precision [KH28, Mar29a]. Predominating [WM29]. Present [Car28a]. Principal [Tow29]. Principles [Osb29]. Printing [Clo27]. Probabilities [Mol22]. Probability [Cam23b, Mol27, Tho26]. Problem [Bla28]. Problems [Car28a, Cha25, Mol22, Mol27]. Process [Har28a]. Product [Dod28]. Production [GHM27, Pet28]. Projector [Scr29]. Propagation [Car26b, Car26c, CH27, Mea25, NS25]. Properties [Elm29b, TS29, Tow29]. Public [GM23, MC23]. Purified [GW29].

Quality [Her29, MH26, She26c, She27]. Quanta [Dar25b, Dav29]. Quartz [Lac29].

Radiation [Dar29c, Wis29]. Radio [And26, AE23b, BDW29, BMP26, Bro27, Esp22, EAB25, Esp27, Har23, Nel27, NE23]. Radioactivity [Dar27a]. Random [MC24]. Range [MC27]. Rating [Dod28]. Ray [Joh22]. Reactance [Fos24a]. Reactances [Car25c]. Recapitulation [Dar26b]. Receiver [WT28]. Receivers [Cur25]. Receiving [BDW29]. Recently [Smi24]. Reciprocal [Car24b]. Recorder [Fri26]. Recorders [Whe24]. Recording [Fre29, Mac29, Mar29b, MH26]. Recurrent [Mas27, Zob28]. Reference [Mar24, MG29]. Reflection [DG29]. Regarding [Fos24b]. Regular [Mas27]. Regulation [Wil26]. Relation [Hel22, Tru22]. Relations [Har23]. Relays [Mil24, SP24]. Rental [Hel22]. Rents [Hel22]. Repeaters [Cri25b]. Representation [Car25c]. Reproducing [MH26]. Research [Cam24, MH26]. Resistance [FB28, Mea25, Zob28]. Resistances [Car25c]. Results [Cog28]. Return [Car26c, Car29]. Review [Tow29]. Rigorous [Car28b].

Sampled [MW29]. Samples [MC24, She26a]. Sampling [Cog28, DR29]. saving [Kir23c]. Scattering [Dar29a]. Sea [NE23]. Second [Dar25d, Dar28b]. Section [Pil22]. Selective [Car25b]. Service [Cra28, Wat28]. Sheet [Hoc26, TG29, TS29]. Shielded [Fer27, Sha27]. Shielding [Fer29]. Ships [NE23]. Showing [Cam23b]. Side [Har23]. Side-Bands [Har23]. Signal [Her29]. Signaling [And26]. Signals [GHM27]. Significance [MC24]. Simulating [Hoy23, Hoy24]. Sinusoidal [Car24a]. Sleet [Kir23a]. Small [Her28, She26a]. Smooth [Hoy23]. Solids [Dar24b]. Solutions [Mas29]. Some [BMP26, Car28a, Cog28, Dar23, Dar24c, Dar24a, Dar24d, Dar24b, Dar25a, Dar25b, Dar25c, Dar25d, Dar26b, Nan24, She24, TS29]. Sound [Mac29, Scr29, Sto29]. Sounds [CS24, Cra25, Cra27, HF22, SB26]. Source [Fer24]. Sources [Ano26c, Ano26d, Ano26b, Ano27d, Ano29b, Ano29c, Ano29d]. Speaker [Bos29, WT28]. Specializing [Kir23b]. Speech [CM22, Cra25, Fle22, Fle25, MH26, Sac25, SB26, Sha27, Siv29]. Speed [Buc28, Nyq24, Sto29]. Springs [Tow29]. Standard [Mar29a]. State [Bos29]. Static [Car25b, Fri26]. Statistical [Dar29c, She24]. Status [Car28a]. Steady [Bos29]. Stethophone [FD24].

Stethoscope [FD24]. Storm [Kir23a]. Studies [BMP26, Rho25, TG29]. Study [CS24, Cra27, Ger29, Mas27]. Submarine [Buc25, CG22, Clo27, Cur27, Gil26, Hit26]. Summation [Cam23b, Mol29, Tho26]. Switchboard [Dav27]. Switching [CMC23]. Synchronization [SM27b, Sto29]. Synchronized [Sto29]. System [And26, Ano24a, Ano25a, BDW29, CH27, CMC23, GG27, MC23, MG29, Nel27, NE23, Scr29, Tru22, Ano24a, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano26c, Ano26d, Ano26b, Ano27a, Ano27b, Ano27c, Ano27d, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano29b, Ano29c, Ano29d, Cha25, Kir23a, Nan24, Ros23, SP24]. Systems [ADG28, GM23, Mar24].

Tank [Smi25]. Tapered [Mas27]. Technical [Ano24a, Ano24b, Ano25a, Ano25b, Ano25c, Ano25d, Ano26a, Ano26c, Ano26d, Ano26b, Ano27a, Ano27b, Ano27c, Ano27d, Ano28a, Ano28b, Ano28c, Ano28d, Ano29a, Ano29b, Ano29c, Ano29d, Bla28, Luc24]. Telegraph [Buc25, Clo27, Cur27, Gil26, Gil27, Her29, Nyq24, Ros23, Sha26, Wil26]. Telegraphy [Buc28]. Telephone [ADG28, Cla23, CMC23, Cri25a, Cri25b, Cur25, Dem23, EH27, EAB25, Fle26, GW29, Har24, Har25, Har28a, IHPC25, KH28, Kir23b, Luc24, MC23, Mar24, MG29, MH26, Mil24, Mol22, Mol27, Nan24, NE23, Ros23, SF26, Tow29, WS27, You27]. Telephony [AE23b, BDW29, Bla28, Bro27, Wat28, Wol25]. Television [GG27, GHM27, Ive27, Nel27, SM27b]. Temperature [Lac29]. Terminal [Car26a]. Terms [FB28]. Test [DG29, TS29]. Testing [FS29]. Tests [Har24]. Textile [GW29]. Textiles [WM29]. Theaters [Scr29]. Their [Fle23, Har24, Kin23]. Theorem [Car24b, Fos24a, Pen29]. Theorems [Fos24b]. Theories [Car28b, Dar29c]. Theory [Cam22b, Car25a, Car26b, Car26a, Car27, Car28a, Cog28, Dar28a, Dar28b, Fle23, Fle26, Mol22, Mol27, Zob23]. Thermionic [Kin23, Lle26]. These [Tow29]. Third [Dar26b]. Time [Her28]. Toll [Dav27, EH27]. Tool [KH28]. Traffic [MC24]. Transatlantic [AE23b, BDW29, Bla28, Bro27, EAB25, Wat28]. Transient [CZ23, Mar29b, Mas29]. Transmission [BMP26, CG22, Car28a, Car28b, Cla23, Esp22, EAB25, GG27, Har24, Har25, Har23, Har28b, Her29, IHPC25, Mar24, Mar29c, MG29, Nel27, Smi24, Zob24]. Transportation [Kir23b]. Treating [Smi25]. Trunking [Mol22, Mol27]. Tube [Cur27, Hor24, Joh29, Lle26, Wil22]. Tubes [Kin23, PE27]. Tubular [Mea25]. Tuning [Fer24]. Two [Car25c, Cri25b, Fos24b]. Two-Mesh [Fos24b]. Two-Way [Cri25b]. Type [WT28, Wil22]. Types [Tow29].

Underground [Car29]. Uniform [Zob23]. Unit [Mar24, Mar29c, Smi24]. Universe [MW29]. Unknown [MW29]. Upon [WM29]. Use [Cog28, Kir23c, MC23, Scr29]. Used [PE27]. Useful [Fle25]. Utilization [GHM27].

V [Dar24b]. Vacuum

[Cur27, Hor24, Kin23, Lle26, PE27, Wil22]. Values [Hel22]. Valve [Mac29]. Very [AE23a, Dar26b, Nan24]. VI [Car26b, Dar25a]. Vibration [Lac29]. Vibratory [Cur25]. VII [Dar25b]. VIII [Dar25c]. Voice [MC27]. Voltage [Car26b, Joh22, Wol25]. Vowel [CS24, Cra27].

Wave [BDW29, Cam22b, CZ23, Car26c, Dar27c, JS25, Mea25, Zob23, Zob24]. Wave-Filter [Cam22b]. Wave-Filters [Zob24, CZ23, Zob23]. Waves [Dar25b, Dar29b, DG29, Lan27, NS25]. Wax [Fre29]. Way [Cri25b]. Wire [Car28a, Car29, Esp22, GG27, SM27a].

Wires [Car26c, CH27, Car28b]. Wiring [GW29]. Work [Har25, Kir23b, Kir23c].

X [Dar26b]. XI [Dar26a]. XII [Dar27a]. XIII [Dar27b]. XIV [Dar27c]. XV [Dar28a]. XVI [Dar28b]. XVII [Dar29a]. XVIII [Dar29b].

York [And26, Pil22, WS27].

References

Affel:1928:CSL

[ADG28] H. A. Affel, C. S. Demarest, and C. W. Green. Carrier systems on long distance telephone lines. The Bell System Technical Journal, 7(3):564–629, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-564.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-564.pdf.

Arnold:1923:PNM

[AE23a] H. D. Arnold and G. W. Elmen. Permalloy, A new magnetic material of very high permeability. The Bell System Technical Journal, 2(3):101–111, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-101.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-3-101.pdf.

Arnold:1923:TRT

[AE23b] H. D. Arnold and Lloyd Espenschied. Transatlantic radio telephony. The Bell System Technical Journal, 2(4):116–144, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
4-116.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-4-116.pdf.

Anderson:1926:RSS

[And26] S. E. Anderson. Radio signaling system for the New York Police Department. The Bell System Technical Journal, 5(4):529–538, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-529.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-529.pdf.

Anonymous:1922:CIa

[Ano22a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 1(1):145–146, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
1-145.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-1-145.pdf.

Anonymous:1922:CIb

[Ano22b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 1(2):152–153, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-

2-152.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-152.pdf.

Anonymous:1922:F

[Ano22c] Anonymous. Forward. The Bell System Technical Journal, 1(1):1–3, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-1-1.pdf.

Anonymous:1923:CIa

[Ano23a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 2(1):133, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
1-133.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-1-133.pdf.

Anonymous:1923:CIb

[Ano23b] Anonymous. The contributors to this issue. The Bell System Technical Journal, 2(2):162–163, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-162.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-162.pdf.

Anonymous:1923:CIc

[Ano23c] Anonymous. The contributors to this issue. The Bell System Technical Journal, 2(3):186–187, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-186.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-3-186.pdf.

Anonymous:1923:CId

[Ano23d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 2(4):181–182, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
4-181.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-4-181.pdf.

Anonymous:1924:ABS

[Ano24a] Anonymous. Abstracts of Bell System technical papers not appearing in the Bell System Technical Journal. The Bell System Technical Journal, 3(2):347–350, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-347.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-347.pdf.

Anonymous:1924:ATP

[Ano24b] Anonymous. Abstracts of technical papers. The Bell System Technical Journal, 3(3):525–528, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-525.pdf;   http://www.alcatel-

lucent.com/bstj/vol03-1924/articles/
bstj3-3-525.pdf.

Anonymous:1924:CIa

[Ano24c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 3(1):179–180, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
1-179.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-1-179.pdf.

Anonymous:1924:CIb

[Ano24d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 3(2):351–352, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-351.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-351.pdf.

Anonymous:1924:CIc

[Ano24e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 3(3):529–530, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-529.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-529.pdf.

Anonymous:1924:CId

[Ano24f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 3(4):686–687, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
4-686.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-4-686.pdf.

Anonymous:1925:ABSa

[Ano25a] Anonymous. Abstracts of Bell System technical papers not appearing in the Bell System Technical Journal. The Bell System Technical Journal, 4(1):178–183, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-178.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-178.pdf.

Anonymous:1925:ABSb

[Ano25b] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 4(2):339–346, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
2-339.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-2-339.pdf.

Anonymous:1925:ABSc

[Ano25c] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 4(3):508–511, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol04/bstj4-
3-508.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-3-508.pdf.

Anonymous:1925:ABSd

[Ano25d] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 4(4):762–765, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-762.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-762.pdf.

Anonymous:1925:CIa

[Ano25e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 4(1):184–185, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-184.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-184.pdf.

Anonymous:1925:CIb

[Ano25f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 4(2):347–348, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
2-347.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-2-347.pdf.

Anonymous:1925:CIc

[Ano25g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 4(3):512–513, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-512.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-512.pdf.

Anonymous:1925:CId

[Ano25h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 4(4):766–767, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        4-766.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-4-766.pdf.

Anonymous:1926:ABS

[Ano26a] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 5 (1):214–218, January 1926. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        1-214.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-1-214.pdf.

Anonymous:1926:ARTc

[Ano26b] Anonymous. Abstracts of recent technical books and papers from Bell System sources. The Bell System Technical Journal, 5(4):652–659, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-652.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-652.pdf.

Anonymous:1926:ARTa

[Ano26c] Anonymous. Abstracts of recent technical papers from Bell System sources. The Bell System Technical Journal, 5(2):385–389, April 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
2-385.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-2-385.pdf.

Anonymous:1926:ARTb

[Ano26d] Anonymous. Abstracts of recent technical papers from Bell System sources. The Bell System Technical Journal, 5(3):524–525, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
3-524.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-3-524.pdf.

Anonymous:1926:CIa

[Ano26e] Anonymous. Contributors in this issue. The Bell System Technical Journal, 5(1):219–220, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-

                                        1-219.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-1-219.pdf.

Anonymous:1926:CIb

[Ano26f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 5(2):390–391, April 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        2-390.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-2-390.pdf.

Anonymous:1926:CIc

[Ano26g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 5(3):526–527, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        3-526.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-3-526.pdf.

Anonymous:1926:CId

[Ano26h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 5(4):660–661, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        4-660.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-4-660.pdf.

Anonymous:1927:ABSa

[Ano27a] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 6(2):367–371, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
2-367.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-2-367.pdf.

Anonymous:1927:ABSb

[Ano27b] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 6(3):546–548, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-546.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-546.pdf.

Anonymous:1927:ABSc

[Ano27c] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 6(4):750–751, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-750.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-750.pdf.

Anonymous:1927:ART

[Ano27d] Anonymous. Abstracts of recent technical papers from Bell System sources. The Bell System Technical Journal, 6(1):181–183, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        1-181.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-1-181.pdf.

Anonymous:1927:CIa

[Ano27e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 6(1):184–185, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        1-184.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-1-184.pdf.

Anonymous:1927:CIb

[Ano27f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 6(2):372–373, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        2-372.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-2-372.pdf.

Anonymous:1927:CIc

[Ano27g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 6(3):549–550, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        3-549.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-3-549.pdf.

Anonymous:1927:CId

[Ano27h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 6(4):752–754, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-752.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-752.pdf.

Anonymous:1927:LE

[Ano27i] Anonymous. Letters to the editor. The Bell System Technical Journal, 6(1):172–180, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-172.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-172.pdf.

Anonymous:1928:ABSa

[Ano28a] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 7 (1):154–158, January 1928. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
1-154.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-1-154.pdf.

Anonymous:1928:ABSb

[Ano28b] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 7(2):369–372, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-369.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-369.pdf.

Anonymous:1928:ABSc

[Ano28c] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 7(3):630–635, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-630.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-630.pdf.

Anonymous:1928:ABSd

[Ano28d] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 7(4):808–813, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
4-808.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-4-808.pdf.

Anonymous:1928:CIa

[Ano28e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 7(1):159–160, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
1-159.pdf;   http://www.alcatel-

lucent.com/bstj/vol07-1928/articles/
bstj7-1-159.pdf.

Anonymous:1928:CIb

[Ano28f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 7(2):373–374, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-373.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-373.pdf.

Anonymous:1928:CIc

[Ano28g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 7(3):636–637, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-636.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-636.pdf.

Anonymous:1928:CId

[Ano28h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 7(4):814, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
4-814.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-4-814.pdf.

Anonymous:1928:JMI

[Ano28i] Anonymous. Joint meeting of the Institution of Electrical Engineers and the American Institute of Electrical Engineers. The Bell System Technical Journal, 7(2):161–167, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-161.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-161.pdf.

Anonymous:1929:ABS

[Ano29a] Anonymous. Abstracts of Bell System technical papers not appearing in this journal. The Bell System Technical Journal, 8 (1):209–213, January 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-209.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-209.pdf.

Anonymous:1929:ATAa

[Ano29b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 8(2):429–432, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-429.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-429.pdf.

Anonymous:1929:ATAb

[Ano29c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 8(3):605–610, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-605.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-605.pdf.

Anonymous:1929:ATAc

[Ano29d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 8(4):855–859, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-855.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-4-855.pdf.

Anonymous:1929:CIb

[Ano29e] Anonymous. Contributions to this issue. The Bell System Technical Journal, 8(2):433–434, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-433.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-433.pdf.

Anonymous:1929:CIc

[Ano29f] Anonymous. Contributions to this issue. The Bell System Technical Journal, 8(3):611–612, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-611.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-611.pdf.

Anonymous:1929:CIa

[Ano29g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 8(1):214–216, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        1-214.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-1-214.pdf.

Anonymous:1929:CId

[Ano29h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 8(4):860–861, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        4-860.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-4-860.pdf.

Bateman:1923:MGA

[Bat23] Helene C. Bateman. A method of graphical analysis. The Bell System Technical Journal, 2(3):77–100, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        3-77.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-3-77.pdf.

Bailey:1929:RSL

[BDW29] Austin Bailey, S. W. Dean, and W. T. Wintringham. The receiving system for long-wave transatlantic radio telephony. The Bell System Technical Journal, 8(2):309–367, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-309.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-309.pdf.

Blackwell:1928:TTT

[Bla28] O. B. Blackwell. Transatlantic telephony — the technical problem. The Bell System Technical Journal, 7(2):168–186, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-168.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-168.pdf.

Bown:1926:SSR

[BMP26] Ralph Bown, DeLoss K. Martin, and Ralph K. Potter. Some studies in radio broadcast transmission. The Bell System Technical Journal, 5(1):143–213, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
1-143.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-1-143.pdf.

Bostwick:1929:ACI

[Bos29] L. G. Bostwick. Acoustic considerations involved in steady state loud speaker measurements. The Bell System Technical Journal, 8 (1):135–158, January 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        1-135.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-1-135.pdf.

Brown:1927:TRT

[Bro27] Ralph Brown. Transatlantic radio telephony. The Bell System Technical Journal, 6(2):248–257, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        2-248.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-2-248.pdf.

Buckley:1925:LST

[Buc25] Oliver E. Buckley. The loaded submarine telegraph cable. The Bell System Technical Journal, 4 (3):355–374, July 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-355.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-355.pdf.

Buckley:1928:HSO

[Buc28] Oliver E. Buckley. High-speed ocean cable telegraphy. The Bell System Technical Journal, 7(2):225–267, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        2-225.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-2-225.pdf.

Campbell:1922:DCM

[Cam22a] George A. Campbell. Direct capacity measurement. The Bell System Technical Journal, 1(1):18–38, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
1-18.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-1-18.pdf.

Campbell:1922:PTE

[Cam22b] George A. Campbell. Physical theory of the electric wavefilter. The Bell System Technical Journal, 1(2):1–32, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-1.pdf.

Campbell:1923:MIG

[Cam23a] George A. Campbell. Mutual impedances of grounded circuits. The Bell System Technical Journal, 2(4):1–30, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
4-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-4-1.pdf.

Campbell:1923:PCS

[Cam23b] George A. Campbell. Probability curves showing Poisson’s exponential summation. The Bell System Technical Journal, 2(1):95–113, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
1-95.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-1-95.pdf.

Campbell:1924:MIR

[Cam24] George A. Campbell. Mathematics in industrial research. The Bell System Technical Journal, 3 (4):550–557, October 1924. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
4-550.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-4-550.pdf.

Campbell:1928:PAF

[Cam28] George A. Campbell. The practical application of the Fourier integral. The Bell System Technical Journal, 7(4):639–707, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
4-639.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-4-639.pdf.

Carson:1922:HOC

[Car22] John R. Carson. The Heaviside operational calculus. The Bell System Technical Journal, 1 (2):43–55, November 1922. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol01/bstj1-
2-43.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-43.pdf.

Carson:1924:BSC

[Car24a] John R. Carson. The building-up of sinusoidal currents in long periodically loaded lines. The Bell System Technical Journal, 3(4):558–566, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
4-558.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-4-558.pdf.

Carson:1924:GRT

[Car24b] John R. Carson. A generalization of the reciprocal theorem. The Bell System Technical Journal, 3(3):393–399, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-393.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-393.pdf.

Carson:1925:ECT

[Car25a] John R. Carson. Electric circuit theory and the operational calculus. The Bell System Technical Journal, 4(4):685–761, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-685.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-685.pdf.

Carson:1925:SCS

[Car25b] John R. Carson. Selective circuits and static interference. The Bell System Technical Journal, 4(2):265–279, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        2-265.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-2-265.pdf.

Carter:1925:GRI

[Car25c] Charles W. Carter, Jr. Graphic representation of the impedance of networks containing resistances and two reactances. The Bell System Technical Journal, 4(3):387–401, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-387.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-387.pdf.

Carson:1926:ECTb

[Car26a] John R. Carson. Electric circuit theory and the operational calculus chapter IX, the finite line with terminal impedances. The Bell System Technical Journal, 5(2):336–384, April 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        2-336.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-2-336.pdf.

Carson:1926:ECTa

[Car26b] John R. Carson. Electric circuit theory and the operational calculus chapter VI, propagation of current and voltage along the noninductive cable. The Bell System Technical Journal, 5(1):50–95, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
1-50.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-1-50.pdf.

Carson:1926:WPO

[Car26c] John R. Carson. Wave propagation in overhead wires with ground return. The Bell System Technical Journal, 5(4):539–554, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-539.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-539.pdf.

Carson:1927:ETF

[Car27] John R. Carson. Electromagnetic theory and the foundations of electric circuit theory. The Bell System Technical Journal, 6(1):1–17, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-1.pdf.

Carson:1928:PSW

[Car28a] John R. Carson. The present status of wire transmission theory and some of its outstanding problems. The Bell System Technical Journal, 7(2):268–280, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        2-268.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-2-268.pdf.

Carson:1928:RAT

[Car28b] John R. Carson. The rigorous and approximate theories of electrical transmission along wires. The Bell System Technical Journal, 7(1):11–25, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        1-11.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-1-11.pdf.

Carson:1929:GRI

[Car29] John R. Carson. Ground return impedance: Underground wire with Earth return. The Bell System Technical Journal, 8(1):94–98, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        1-94.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-1-94.pdf.

Carson:1922:TCS

[CG22] John R. Carson and J. J. Gilbert. Transmission characteristics of the submarine cable. The Bell System Technical Journal, 1(1):88–115, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
1-88.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-1-88.pdf.

Carson:1927:PPC

[CH27] John R. Carson and Ray S.

Hoyt. Propagation of periodic currents over a system of parallel wires. The Bell System Technical Journal, 6(3):495–545, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-495.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-495.pdf.

Charlesworth:1925:GEP

[Cha25] H. P. Charlesworth. General engineering problems of the Bell System. The Bell System Technical Journal, 4(4):515–541, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-515.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-515.pdf.

Clark:1923:TTL

[Cla23] A. B. Clark. Telephone transmission over long cable circuits. The Bell System Technical Journal, 2(1):67–94, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
1-67.pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-1-67.pdf.

Clokey:1927:APE

[Clo27] A. A. Clokey. Automatic printing equipment for long loaded submarine telegraph cables. The Bell System Technical Journal, 6(3):402–424, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        3-402.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-3-402.pdf.

Crandall:1922:AED

[CM22] I. B. Crandall and D. MacKenzie. Analysis of the energy distribution in speech. The Bell System Technical Journal, 1(1):116–128, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol01/bstj1-
                                        1-116.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol01-1922/articles/
                                        bstj1-1-116.pdf.

Craft:1923:MST

[CMC23] E. B. Craft, L. F. Morehouse, and H. P. Charlesworth. Machine switching telephone system for large metropolitan areas. The Bell System Technical Journal, 2(2):53–89, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        2-53.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-2-53.pdf.

Coggins:1928:SGR

[Cog28] Paul P. Coggins. Some general results of elementary sampling theory for engineering use. The Bell System Technical Journal, 7(1):26–69, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
1-26.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-1-26.pdf.

Crandall:1925:SS

[Cra25] Irving B. Crandall. The sounds of speech. The Bell System Technical Journal, 4(4):586–626, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-586.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-586.pdf.

Crandall:1927:DSV

[Cra27] Irving B. Crandall. Dynamical study of the vowel sounds, Part II. The Bell System Technical Journal, 6(1):100–116, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-100.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-100.pdf.

Craft:1928:ACS

[Cra28] Edward B. Craft. Airways communication service. The Bell System Technical Journal, 7(4):797–807, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
4-797.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-4-797.pdf.

Crisson:1925:ILT

[Cri25a] George Crisson. Irregularities in loaded telephone circuits. The Bell System Technical Journal, 4 (4):561–585, October 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-561.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-561.pdf.

Crisson:1925:LGT

[Cri25b] George Crisson. The limitation of the gain of two-way telephone repeaters by impedance irregularities. The Bell System Technical Journal, 4(1):15–25, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-15.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-15.pdf.

Crandall:1924:DSV

[CS24] I. B. Crandall and C. F. Sacia. A dynamical study of the vowel sounds. The Bell System Technical Journal, 3(2):232–237, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-

2-232.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-232.pdf.

Curtis:1925:VCI

[Cur25] A. S. Curtis. The vibratory characteristics and impedance of telephone receivers at low power inputs. The Bell System Technical Journal, 4(3):402–406, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
3-402.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-3-402.pdf.

Curtis:1927:AVT

[Cur27] Austen M. Curtis. The application of vacuum tube amplifiers to submarine telegraph cables. The Bell System Technical Journal, 6(3):425–441, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-425.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-425.pdf.

Carson:1923:TOE

[CZ23] John R. Carson and Otto J. Zobel. Transient oscillations in electric wave-filters. The Bell System Technical Journal, 2(3):1–52, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-3-1.pdf.

Darrow:1923:SCA

[Dar23] K. K. Darrow. Some contemporary advances in physics. The Bell System Technical Journal, 2(4):101–115, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        4-101.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-4-101.pdf.

Darrow:1924:SCAb

[Dar24a] Karl K. Darrow. Some contemporary advances in physics — III. The Bell System Technical Journal, 3(2):268–298, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        2-268.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-2-268.pdf.

Darrow:1924:SCAd

[Dar24b] Karl K. Darrow. Some contemporary advances in physics — V electricity in solids. The Bell System Technical Journal, 3(4):621–650, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        4-621.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-4-621.pdf.

Darrow:1924:SCAa

[Dar24c] Karl K. Darrow. Some contemporary advances in physics-II. The Bell System Technical Journal, 3(1):158–178, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
1-158.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-1-158.pdf.

Darrow:1924:SCAc

[Dar24d] Karl K. Darrow. Some contemporary advances in physics-IV. The Bell System Technical Journal, 3(3):468–494, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-468.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-468.pdf.

Darrow:1925:SCAa

[Dar25a] Karl K. Darrow. Some contemporary advances in physics — VI electricity in gases. The Bell System Technical Journal, 4(1): 112–151, January 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-112.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-112.pdf.

Darrow:1925:SCAb

[Dar25b] Karl K. Darrow. Some contemporary advances in physics — VII waves and quanta. The Bell System Technical Journal, 4(2):280–326, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        2-280.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-2-280.pdf.

Darrow:1925:SCAc

[Dar25c] Karl K. Darrow. Some contemporary advances in physics — VIII the atom-model, first part. The Bell System Technical Journal, 4(3):407–458, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-407.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-407.pdf.

Darrow:1925:SCAd

[Dar25d] Karl K. Darrow. Some contemporary advances in physics IX the atom-model, second part. The Bell System Technical Journal, 4 (4):642–684, October 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        4-642.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-4-642.pdf.

Darrow:1926:CAP

[Dar26a] Karl K. Darrow. Contemporary advances in physics — XI — ionization. The Bell System Technical Journal, 5(3):463–492, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        3-463.pdf;   http://www.alcatel-

lucent.com/bstj/vol05-1926/articles/
bstj5-3-463.pdf.

Darrow:1926:SCA

[Dar26b] Karl K. Darrow. Some contemporary advances in physics-X, the atom-model, third part, A very brief recapitulation of what has gone before. The Bell System Technical Journal, 5(1):96–142, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
1-96.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-1-96.pdf.

Darrow:1927:CAPa

[Dar27a] Karl K. Darrow. Contemporary advances in physics — XII radioactivity. The Bell System Technical Journal, 6(1):55–99, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-55.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-55.pdf.

Darrow:1927:CAPb

[Dar27b] Karl K. Darrow. Contemporary advances in physics-XIII. ferromagnetism. The Bell System Technical Journal, 6(2):295–366, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
2-295.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-2-295.pdf.

Darrow:1927:CAPc

[Dar27c] Karl K. Darrow. Contemporary advances in physics. XIV introduction to wave mechanics. The Bell System Technical Journal, 6(4):653–701, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        4-653.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-4-653.pdf.

Darrow:1928:CAPa

[Dar28a] Karl K. Darrow. Contemporary advances in physics — XV the classical theory of light, first part. The Bell System Technical Journal, 7(2):281–320, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        2-281.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-2-281.pdf.

Darrow:1928:CAPb

[Dar28b] Karl K. Darrow. Contemporary advances in physics, XVI the classical theory of light, second part. The Bell System Technical Journal, 7(4):730–761, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        4-730.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-4-730.pdf.

Darrow:1929:CAPa

[Dar29a] Karl K. Darrow. Contemporary advances in physics, XVII, the scattering of light with change of frequency. The Bell System Technical Journal, 8(1):64–93, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-64.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-64.pdf.

Darrow:1929:CAPb

[Dar29b] Karl K. Darrow. Contemporary advances in physics, XVIII, the diffraction of waves by crystals. The Bell System Technical Journal, 8(2):391–428, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-391.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-391.pdf.

Darrow:1929:STM

[Dar29c] Karl K. Darrow. Statistical theories of matter, radiation and electricity. The Bell System Technical Journal, 8(4):672–748, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-672.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-4-672.pdf.

Davidson:1927:TSN

[Dav27] J. Davidson. Toll switchboard no. 3. The Bell System Technical Journal, 6(1):18–26, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        1-18.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-1-18.pdf.

Davisson:1928:DEC

[Dav28] C. J. Davisson. The diffraction of electrons by a crystal of nickel. The Bell System Technical Journal, 7(1):90–105, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        1-90.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-1-90.pdf.

Davisson:1929:EQ

[Dav29] C. J. Davisson. Electrons and quanta. The Bell System Technical Journal, 8(2):217–224, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        2-217.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-2-217.pdf.

Demarest:1923:TEL

[Dem23] Charles S. Demarest. Telephone equipment for long cable circuits. The Bell System Technical Journal, 2(3):112–140, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-112.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-3-112.pdf.

Davisson:1929:TPE

[DG29] C. J. Davisson and L. H. Germer.

A test for polarization of electron waves by reflection. The Bell System Technical Journal, 8(3):466–481, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-466.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-466.pdf.

Dodge:1928:MRM

[Dod28] H. F. Dodge. A method of rating manufactured product. The Bell System Technical Journal, 7 (2):350–368, April 1928. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-350.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-350.pdf.

Dodge:1929:MSI

[DR29] H. F. Dodge and H. G. Romig.

A method of sampling inspection. The Bell System Technical Journal, 8(4):613–631, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-

                                        4-613.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-4-613.pdf.

Espenschied:1925:TRT

[EAB25] Lloyd Espenschied, C. N. Anderson, and Austin Bailey. Transatlantic radio telephone transmission. The Bell System Technical Journal, 4(3):459–507, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-459.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-459.pdf.

Edwards:1927:LOT

[EH27] P. G. Edwards and H. W. Herrington. The location of opens in toll telephone cables. The Bell System Technical Journal, 6(1):27–54, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        1-27.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-1-27.pdf.

Elmen:1929:MAI

[Elm29a] G. W. Elmen. Magnetic alloys of iron, nickel, and cobalt. The Bell System Technical Journal, 8(3):435–465, July 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        3-435.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-3-435.pdf.

Elmen:1929:MPP

[Elm29b] G. W. Elmen. Magnetic properties of perminvar. The Bell System Technical Journal, 8(1):21–40, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-21.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-21.pdf.

Englund:1928:NPL

[Eng28] C. R. Englund. The natural period of linear conductors. The Bell System Technical Journal, 7(3):404–419, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-404.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-404.pdf.

Espenschied:1922:ARW

[Esp22] Lloyd Espenschied. Application to radio of wire transmission engineering. The Bell System Technical Journal, 1(2):117–141, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-117.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-117.pdf.

Espenschied:1927:RBC

[Esp27] Lloyd Espenschied. Radio broadcast coverage of city areas. The Bell System Technical Journal, 6 (1):117–141, January 1927. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-117.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-117.pdf.

Ferguson:1928:MCT

[FB28] J. G. Ferguson and B. W. Bartlett. The measurement of capacitance in terms of resistance and frequency. The Bell System Technical Journal, 7(3):420–437, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-420.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-420.pdf.

Frederick:1924:QSQ

[FD24] H. A. Frederick and H. F. Dodge. ‘the stethophone,’ an electrical stethoscope. The Bell System Technical Journal, 3(4):531–549, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
4-531.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-4-531.pdf.

Ferguson:1924:CCT

[Fer24] J. G. Ferguson. A clock-controlled tuning fork as a source of constant frequency. The Bell System Technical Journal, 3(1):145–157, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol03/bstj3-
1-145.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-1-145.pdf.

Ferguson:1927:MIS

[Fer27] J. G. Ferguson. Measurement of inductance by the shielded owen bridge. The Bell System Technical Journal, 6(3):375–386, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-375.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-375.pdf.

Ferguson:1929:SHF

[Fer29] J. G. Ferguson. Shielding in high-frequency measurements. The Bell System Technical Journal, 8 (3):560–575, July 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-560.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-560.pdf.

Friis:1924:HFA

[FJ24] H. T. Friis and A. G. Jensen. High frequency amplifiers. The Bell System Technical Journal, 3(2):181–205, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-181.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-181.pdf.

Fletcher:1922:NSI

[Fle22] Harvey Fletcher. The nature of speech and its interpretation. The Bell System Technical Journal, 1(1):129–144, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol01/bstj1-
                                        1-129.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol01-1922/articles/
                                        bstj1-1-129.pdf.

Fletcher:1923:PMA

[Fle23] Harvey Fletcher. Physical measurements of audition and their bearing on the theory of hearing. The Bell System Technical Journal, 2(4):145–180, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        4-145.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-4-145.pdf.

Fletcher:1925:UNC

[Fle25] Harvey Fletcher. Useful numerical constants of speech and hearing. The Bell System Technical Journal, 4(3):375–386, July 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        3-375.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-3-375.pdf.

Fletcher:1926:TOH

[Fle26] Harvey Fletcher. The theory of the operation of the howling telephone with experimental confirmation. The Bell System Technical Journal, 5(1):27–49, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
1-27.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-1-27.pdf.

Foster:1924:RT

[Fos24a] Ronald M. Foster. A reactance theorem. The Bell System Technical Journal, 3(2):259–267, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-259.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-259.pdf.

Foster:1924:TRD

[Fos24b] Ronald M. Foster. Theorems regarding the driving-point impedance of two-mesh circuits. The Bell System Technical Journal, 3(4):651–685, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
4-651.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-4-651.pdf.

Foster:1926:DDA

[Fos26] Ronald M. Foster. Directive diagrams of antenna arrays. The Bell System Technical Journal, 5 (2):292–307, April 1926. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        2-292.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-2-292.pdf.

Frederick:1929:RAW

[Fre29] Halsey A. Frederick. Recent advances in wax recording. The Bell System Technical Journal, 8 (1):159–172, January 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        1-159.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-1-159.pdf.

Friis:1926:SR

[Fri26] H. T. Friis. A static recorder. The Bell System Technical Journal, 5(2):282–291, April 1926. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        2-282.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-2-282.pdf.

Fletcher:1929:ATM

[FS29] H. Fletcher and J. C. Steinberg. Articulation testing methods. The Bell System Technical Journal, 8(4):806–854, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        4-806.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-4-806.pdf.

Germer:1929:AED

[Ger29] L. H. Germer. An application of electron diffraction to the study of gas adsorption. The Bell System Technical Journal, 8(3):591–604, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-591.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-591.pdf.

Gannett:1927:WTS

[GG27] D. K. Gannett and E. I. Green.

Wire transmission system for television. The Bell System Technical Journal, 6(4):616–632, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-616.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-616.pdf.

Gray:1927:PUT

[GHM27] Frank Gray, J. W. Horton, and R. C. Mathes. The production on utilization of television signals. The Bell System Technical Journal, 6(4):560–604, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-560.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-560.pdf.

Gill:1925:OH

[Gil25] F. Gill. Oliver heaviside. The Bell System Technical Journal, 4 (3):349–354, July 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
3-349.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-3-349.pdf.

Gilbert:1926:EIS

[Gil26] J. J. Gilbert. Extraneous interference on submarine telegraph cables. The Bell System Technical Journal, 5(3):404–417, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
3-404.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-3-404.pdf.

Gilbert:1927:DEC

[Gil27] J. J. Gilbert. Determination of electrical characteristics of loaded telegraph cables. The Bell System Technical Journal, 6(3):387–401, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-387.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-387.pdf.

Gherardi:1926:JHA

[GK26] Bancroft Gherardi and Robert W. King. Joseph Henry, the American pioneer in electrical communication. The Bell System Technical Journal, 5(1):1–10, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-1-1.pdf.

Green:1923:PAS

[GM23] I. W. Green and J. P. Maxfield. Public address systems. The Bell System Technical Journal, 2(2):113–142, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-113.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-113.pdf.

Glenn:1929:PTI

[GW29] H. H. Glenn and E. B. Wood. Purified textile insulation for telephone central office wiring. The Bell System Technical Journal, 8(2):243–256, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-243.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-243.pdf.

Hartley:1923:RCS

[Har23] R. V. L. Hartley. Relations of carrier and side-bands in radio transmission. The Bell System Technical Journal, 2(2):90–112, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-90.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-90.pdf.

Harden:1924:ETT

[Har24] W. H. Harden. Electrical tests and their applications in the maintenance of telephone transmission. The Bell System Technical Journal, 3(3):353–392, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        3-353.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-3-353.pdf.

Harden:1925:PTT

[Har25] W. H. Harden. Practises in telephone transmission maintenance work. The Bell System Technical Journal, 4(1):26–51, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        1-26.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-1-26.pdf.

Hart:1928:RDP

[Har28a] C. D. Hart. Recent developments in the process of manufacturing lead-covered telephone cable. The Bell System Technical Journal, 7(2):321–342, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        2-321.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-2-321.pdf.

Hartley:1928:TI

[Har28b] R. V. L. Hartley. Transmission of information. The Bell System Technical Journal, 7(3):535–563, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-535.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-535.pdf.

Helmle:1922:RBR

[Hel22] W. C. Helmle. The relation between rents and incomes, and the distribution of rental values. The Bell System Technical Journal, 1(2):82–109, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-82.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-82.pdf.

Herman:1928:BMS

[Her28] J. Herman. Bridge for measuring small time intervals. The Bell System Technical Journal, 7(2):343–349, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-343.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-343.pdf.

Herman:1929:ESD

[Her29] J. Herman. Effect of signal distortion on Morse telegraph transmission quality. The Bell System Technical Journal, 8(2):267–285, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-267.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-267.pdf.

Hartley:1922:BLC

[HF22] R. V. L. Hartley and Thornton C. Fry. The binaural location of complex sounds. The Bell System Technical Journal, 1(2):33–42, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-33.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-33.pdf.

Hitchcock:1926:CCC

[Hit26] H. W. Hitchcock. Carrier-current communication on submarine cables. The Bell System Technical Journal, 5(4):636–651, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-636.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-636.pdf.

Hoch:1922:PLI

[Hoc22] E. T. Hoch. Power losses in insulating materials. The Bell System Technical Journal, 1(2):110–116, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-110.pdf;   http://www.alcatel-

lucent.com/bstj/vol01-1922/articles/
bstj1-2-110.pdf.

Hoch:1926:EEM

[Hoc26] E. T. Hoch. Electrode effects in the measurement of power factor and dielectric constant of sheet insulating materials. The Bell System Technical Journal, 5(4):555–572c, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-555.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-555.pdf.

Horton:1924:VTO

[Hor24] J. W. Horton. Vacuum tube oscillators — a graphical method of analysis. The Bell System Technical Journal, 3(3):508–524, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-508.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-508.pdf.

Hoyt:1923:ISL

[Hoy23] Ray S. Hoyt. Impedance of smooth lines, and design of simulating networks. The Bell System Technical Journal, 2(2):1–40, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-1.pdf.

Hoyt:1924:ILL

[Hoy24] Ray S. Hoyt. Impedance of loaded lines, and design of simulating and compensating networks. The Bell System Technical Journal, 3(3):414–467, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        3-414.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-3-414.pdf.

Harris:1923:MGE

[HS23] J. E. Harris and E. E. Schumacher. Measurements on the gases evolved from glasses of known chemical composition. The Bell System Technical Journal, 2(1):122–132, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        1-122.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-1-122.pdf.

Ives:1925:TPT

[IHPC25] H. E. Ives, J. W. Horton, R. D. Parker, and A. B. Clark. The transmission of pictures over telephone lines. The Bell System Technical Journal, 4(2):187–214, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        2-187.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-2-187.pdf.

Ives:1926:AMP

[Ive26] Herbert E. Ives. The alkali metal photoelectric cell. The Bell System Technical Journal, 5(2):320–335, April 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
2-320.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-2-320.pdf.

Ives:1927:T

[Ive27] Herbert E. Ives. Television. The Bell System Technical Journal, 6 (4):551–559, October 1927. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-551.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-551.pdf.

Johnson:1922:LVC

[Joh22] J. B. Johnson. A low voltage cathode ray oscillograph. The Bell System Technical Journal, 1 (2):142–151, November 1922. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-142.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-142.pdf.

Johnson:1929:BTH

[Joh29] J. B. Johnson. A braun tube hysteresigraph. The Bell System Technical Journal, 8(2):286–308, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-286.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-286.pdf.

Johnson:1925:MIW

[JS25] K. S. Johnson and T. E. Shea. Mutual inductance in wave filters with an introduction on filter design. The Bell System Technical Journal, 4(1):52–111, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-52.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-52.pdf.

Kasley:1928:PTM

[KH28] J. H. Kasley and F. P. Hutchison. Precision tool making for the manufacture of telephone apparatus. The Bell System Technical Journal, 7(3):375–403, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
3-375.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-3-375.pdf.

King:1923:TVT

[Kin23] Robert W. King. Thermionic vacuum tubes and their applications. The Bell System Technical Journal, 2(4):31–100, October 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol02/bstj2-
4-31.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-4-31.pdf.

Kirk:1923:BSS

[Kir23a] J. N. Kirk. Bell System sleet storm map. The Bell System Technical Journal, 2(1):114–121, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
1-114.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-1-114.pdf.

Kirk:1923:STE

[Kir23b] J. N. Kirk. Specializing transportation equipment in order to adapt it most economically to telephone construction and maintenance work. The Bell System Technical Journal, 2(1):47–66, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
1-47.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-1-47.pdf.

Kirk:1923:ULS

[Kir23c] J. N. Kirk. Use of labor-saving apparatus in outside plant construction work. The Bell System Technical Journal, 2(3):53–76, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-53.pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-3-53.pdf.

Lack:1929:OMV

[Lac29] F. R. Lack. Observations on modes of vibration and temperature coefficients of quartz crystal plates. The Bell System Technical Journal, 8(3):515–535, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol08/bstj8-
                                        3-515.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-3-515.pdf.

Landeen:1927:ACE

[Lan27] A. G. Landeen. Analyzer for complex electric waves. The Bell System Technical Journal, 6(2):230–247, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        2-230.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-2-230.pdf.

Llewellyn:1926:OTV

[Lle26] F. B. Llewellyn. Operation of thermionic vacuum tube circuits. The Bell System Technical Journal, 5(3):433–462, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        3-433.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-3-433.pdf.

Lucas:1924:PTM

[Luc24] Francis F. Lucas. Photomicrography and technical microscopy in its application to telephone apparatus. The Bell System Technical Journal, 3(1):100–144, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
1-100.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-1-100.pdf.

MacKenzie:1929:SRL

[Mac29] Donald MacKenzie. Sound recording with the light valve. The Bell System Technical Journal, 8 (1):173–183, January 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-173.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-173.pdf.

Martin:1924:TUT

[Mar24] W. H. Martin. The transmission unit and telephone transmission reference systems. The Bell System Technical Journal, 3(3):400–408, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-400.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-400.pdf.

Marrison:1929:HPS

[Mar29a] W. A. Marrison. A high precision standard of frequency. The Bell System Technical Journal, 8(3):493–514, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-493.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-493.pdf.

Marrison:1929:ORT

[Mar29b] W. A. Marrison. Oscillographs for recording transient phenomena. The Bell System Technical Journal, 8(2):368–390, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-368.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-368.pdf.

Martin:1929:DNT

[Mar29c] W. H. Martin. Decibel–the name for the transmission unit. The Bell System Technical Journal, 8 (1):1–2, January 1929. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-1.pdf.

Mason:1927:SRC

[Mas27] W. P. Mason. A study of the regular combination of acoustic elements, with applications to recurrent acoustic filters, tapered acoustic fibers, and horns. The Bell System Technical Journal, 6(2):258–294, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
2-258.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-2-258.pdf.

Mason:1929:NMO

[Mas29] W. P. Mason. A new method for obtaining transient solutions of electrical networks. The Bell System Technical Journal, 8(1):109–134, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-109.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-109.pdf.

Martin:1923:UPA

[MC23] W. H. Martin and A. B. Clark. Use of public address system with telephone lines. The Bell System Technical Journal, 2(2):143–161, April 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-143.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-143.pdf.

Molina:1924:DRS

[MC24] E. C. Molina and R. P. Crowell. Deviation of random samples from average conditions and significance to traffic men. The Bell System Technical Journal, 3(1):88–99, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-

                                        1-88.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-1-88.pdf.

Moore:1927:AVF

[MC27] C. R. Moore and A. S. Curtis. An analyzer for the voice frequency range. The Bell System Technical Journal, 6(2):217–229, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        2-217.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-2-217.pdf.

Mead:1925:WPP

[Mea25] Sallie Pero Mead. Wave propagation over parallel tubular conductors: The alternating current resistance. The Bell System Technical Journal, 4(2):327–338, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        2-327.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-2-327.pdf.

Mead:1928:PDP

[Mea28] Sallie Pero Mead. Phase distortion and phase distortion correction. The Bell System Technical Journal, 7(2):195–224, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        2-195.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-2-195.pdf.

Martin:1929:MRS

[MG29] W. H. Martin and C. H. G. Gray.

Master reference system for telephone transmission. The Bell System Technical Journal, 8(3):536–559, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-536.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-536.pdf.

Maxfield:1926:MHQ

[MH26] J. P. Maxfield and H. C. Harrison. Methods of high quality recording and reproducing of music and speech based on telephone research. The Bell System Technical Journal, 5(3):493–523, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
3-493.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-3-493.pdf.

Miller:1924:DCE

[Mil24] D. D. Miller. Design characteristics of electromagnets for telephone relays. The Bell System Technical Journal, 3(2):206–231, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-206.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-206.pdf.

Molina:1922:TPA

[Mol22] Edward C. Molina. The theory of probabilities applied to telephone trunking problems. The Bell System Technical Journal, 1 (2):69–81, November 1922. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
2-69.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-2-69.pdf.

Molina:1927:ATP

[Mol27] Edward C. Molina. Application of the theory of probability to telephone trunking problems. The Bell System Technical Journal, 6(3):461–494, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-461.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-461.pdf.

Molina:1929:ABS

[Mol29] E. C. Molina. Application to the binomial summation of a Laplacian method for the evaluation of definite integrals. The Bell System Technical Journal, 8(1):99–108, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-99.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-99.pdf.

Molina:1929:FDU

[MW29] E. C. Molina and R. I. Wilkinson.

The frequency distribution of the unknown mean of a sampled universe. The Bell System Technical Journal, 8(4):632–645, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-632.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-4-632.pdf.

Nance:1924:SVL

[Nan24] H. H. Nance. Some very long telephone circuits of the Bell System. The Bell System Technical Journal, 3(3):495–507, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
3-495.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-3-495.pdf.

Nichols:1923:RET

[NE23] H. W. Nichols and Lloyd Espenschied. Radio extension of the telephone system to ships at sea. The Bell System Technical Journal, 2(3):141–185, July 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
3-141.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-3-141.pdf.

Nelson:1927:RTS

[Nel27] Edward L. Nelson. Radio transmission system for television. The Bell System Technical Journal, 6 (4):633–652, October 1927. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        4-633.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-4-633.pdf.

Nichols:1925:PEW

[NS25] H. W. Nichols and J. C. Schelleng. Propagation of electric waves over the Earth. The Bell System Technical Journal, 4(2):215–234, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol04/bstj4-
                                        2-215.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol04-1925/articles/
                                        bstj4-2-215.pdf.

Nyquist:1924:CFA

[Nyq24] H. Nyquist. Certain factors affecting telegraph speed. The Bell System Technical Journal, 3(2):324–346, April 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        2-324.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-2-324.pdf.

Osborne:1929:PEC

[Osb29] H. S. Osborne. The principles of electric circuits applied to communication. The Bell System Technical Journal, 8(1):3–20, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-3.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-3.pdf.

Peterson:1927:MVT

[PE27] Eugene Peterson and Herbert P.

Evans. Modulation in vacuum tubes used as amplifiers. The Bell System Technical Journal, 6(3):442–460, July 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
3-442.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-3-442.pdf.

Pennock:1925:EPM

[Pen25] G. A. Pennock. Engineering planning for manufacture. The Bell System Technical Journal, 4(4): 542–560, October 1925. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-542.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-542.pdf.

Pennell:1929:GHE

[Pen29] W. O. Pennell. A generalization of Heaviside’s expansion theorem. The Bell System Technical Journal, 8(3):482–492, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-

                                        3-482.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-3-482.pdf.

Peterson:1928:HPF

[Pet28] Eugene Peterson. Harmonic production in ferromagnetic materials at low frequencies and low flux densities. The Bell System Technical Journal, 7(4):762–796, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        4-762.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-4-762.pdf.

Pilliod:1922:PPS

[Pil22] James J. Pilliod. Philadelphia– Pittsburgh section of the New York– Chicago cable. The Bell System Technical Journal, 1(1):60–87, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol01/bstj1-
                                        1-60.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol01-1922/articles/
                                        bstj1-1-60.pdf.

Peterson:1928:GCM

[PK28] Eugene Peterson and Clyde R. Keith. Grid current modulation. The Bell System Technical Journal, 7(1):106–139, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        1-106.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-1-106.pdf.

Rhodes:1925:ECS

[Rho25] F. L. Rhodes. Engineering cost studies. The Bell System Technical Journal, 4(1):1–14, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-1.pdf.

Robbins:1928:AMG

[Rob28] C. W. Robbins. Automatic machine gaging. The Bell System Technical Journal, 7(4):708–729, October 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
4-708.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-4-708.pdf.

Rose:1923:PAC

[Ros23] Arthur F. Rose. Practical application of carrier telephone and telegraph in the Bell System. The Bell System Technical Journal, 2(2):41–52, April 1923. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol02/bstj2-
2-41.pdf;   http://www.alcatel-
lucent.com/bstj/vol02-1923/articles/
bstj2-2-41.pdf.

Sacia:1925:SPE

[Sac25] C. F. Sacia. Speech power and energy. The Bell System Technical Journal, 4(4):627–641, October 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
4-627.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-4-627.pdf.

Sacia:1926:PFS

[SB26] C. F. Sacia and C. J. Beck. The power of fundamental speech sounds. The Bell System Technical Journal, 5(3):393–403, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
3-393.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-3-393.pdf.

Scriven:1929:SPS

[Scr29] E. O. Scriven. A sound projector system for use in motion picture theaters. The Bell System Technical Journal, 8(1):196–208, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-196.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-196.pdf.

Shaw:1926:DAL

[SF26] Thomas Shaw and William Fondiller. Development and application of loading for telephone circuits. The Bell System Technical Journal, 5 (2):221–281, April 1926. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol05/bstj5-
2-221.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-2-221.pdf.

Shackelton:1928:EMC

[SF28] W. J. Shackelton and J. G. Ferguson.

Electrical measurement of communication apparatus. The Bell System Technical Journal, 7(1):70–89, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
1-70.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-1-70.pdf.

Shanck:1926:NTC

[Sha26] R. B. Shanck. Neutralization of telegraph crossfire. The Bell System Technical Journal, 5(3):418–432, July 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
3-418.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-3-418.pdf.

Shackelton:1927:SBI

[Sha27] W. J. Shackelton. A shielded bridge for inductive impedance measurements at speech and carrier frequencies. The Bell System Technical Journal, 6(1):142–171, January 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
1-142.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-1-142.pdf.

Shewhart:1924:SAS

[She24] W. A. Shewhart. Some applications of statistical methods to the analysis of physical and engineering data. The Bell System Technical Journal, 3(1):43–87, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        1-43.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-1-43.pdf.

Shewhart:1926:CDEb

[She26a] W. A. Shewhart. Correction of data for errors of averages obtained from small samples. The Bell System Technical Journal, 5(2):308–319, April 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        2-308.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-2-308.pdf.

Shewhart:1926:CDEa

[She26b] W. A. Shewhart. Correction of data for errors of measurement. The Bell System Technical Journal, 5(1):11–26, January 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol05/bstj5-
                                        1-11.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol05-1926/articles/
                                        bstj5-1-11.pdf.

Shewhart:1926:QCC

[She26c] W. A. Shewhart. Quality control charts. The Bell System Technical Journal, 5(4):593–603, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-593.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-593.pdf.

Shewhart:1927:QC

[She27] W. A. Shewhart. Quality control. The Bell System Technical Journal, 6(4):722–735, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-722.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-722.pdf.

Siegmund:1929:AEC

[Sie29] H. O. Siegmund. The aluminum electrolytic condenser. The Bell System Technical Journal, 8(1):41–63, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-41.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-41.pdf.

Sivian:1929:SPM

[Siv29] L. J. Sivian. Speech power and its measurement. The Bell System Technical Journal, 8(4):646–661, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-646.pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol08-1929/articles/
                                        bstj8-4-646.pdf.

Shea:1927:DMC

[SM27a] John R. Shea and Samuel McMullan. Developments in the manufacture of copper wire. The Bell System Technical Journal, 6(2):187–216, April 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        2-187.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-2-187.pdf.

Stoller:1927:ST

[SM27b] H. M. Stoller and E. R. Morton. Synchronization of television. The Bell System Technical Journal, 6(4):604–615, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol06/bstj6-
                                        4-604.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol06-1927/articles/
                                        bstj6-4-604.pdf.

Smith:1924:PAR

[Smi24] C. W. Smith. Practical application of the recently adopted transmission unit. The Bell System Technical Journal, 3(3):409–413, July 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        3-409.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-3-409.pdf.

Smith:1925:OTC

[Smi25] T. C. Smith. Open tank creosoting plants for treating chestnut poles. The Bell System Technical Journal, 4(2):235–264, April 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
2-235.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-2-235.pdf.

Shackleton:1924:RBS

[SP24] S. P. Shackleton and H. W. Purcell. Relays in the Bell System. The Bell System Technical Journal, 3(1):1–42, January 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
1-1.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-1-1.pdf.

Stoller:1929:SSC

[Sto29] H. M. Stoller. Synchronization and speed control of synchronized sound pictures. The Bell System Technical Journal, 8(1):184–195, January 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
1-184.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-1-184.pdf.

Townsend:1929:FSN

[TG29] John R. Townsend and Charles H.

Greenall. Fatigue studies of nonferrous sheet metals. The Bell System Technical Journal, 8(3):576–590, July 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
3-576.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-3-576.pdf.

Thorndike:1926:APP

[Tho26] Frances Thorndike. Applications of Poisson’s probability summation. The Bell System Technical Journal, 5(4):604–624, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-604.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-604.pdf.

Townsend:1929:TAS

[Tow29] J. R. Townsend. Telephone apparatus springs: A review of the principal types and the properties desired of these springs. The Bell System Technical Journal, 8(2):257–266, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-257.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-257.pdf.

Trueblood:1922:RPS

[Tru22] H. M. Trueblood. The relation of the Petersen system of grounding power networks to inductive effects in neighboring communication circuits. The Bell System Technical Journal, 1(1):39–59, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol01/bstj1-
1-39.pdf;   http://www.alcatel-
lucent.com/bstj/vol01-1922/articles/
bstj1-1-39.pdf.

Townsend:1929:PPM

[TS29] J. R. Townsend and W. A. Straw.

Physical properties and methods of test for some sheet non-ferrous metals. The Bell System Technical Journal, 8(4):749–805, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-749.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-4-749.pdf.

Waterson:1928:TTS

[Wat28] K. W. Waterson. Transatlantic telephony — service and operating features. The Bell System Technical Journal, 7(2):187–194, April 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
2-187.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-2-187.pdf.

Wente:1928:MAI

[WB28] E. C. Wente and E. H. Bedell. The measurement of acoustic impedance and the absorption coefficient of porous materials. The Bell System Technical Journal, 7(1):1–10, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        1-1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-1-1.pdf.

Wegel:1922:PCA

[Weg22] R. L. Wegel. The physical characteristics of audition and dynamical analysis of the external ear. The Bell System Technical Journal, 1(2):56–68, November 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol01/bstj1-
                                        2-56.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol01-1922/articles/
                                        bstj1-2-56.pdf.

Wheeler:1924:HR

[Whe24] E. B. Wheeler. Humidity recorders. The Bell System Technical Journal, 3(2):238–258, April 1924. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        2-238.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-2-238.pdf.

Wilson:1922:NTH

[Wil22] W. Wilson. A new type of high power vacuum tube. The Bell System Technical Journal, 1(1):4–17, July 1922. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol01/bstj1-
                                        1-4.pdf;   http://www.alcatel-

lucent.com/bstj/vol01-1922/articles/
bstj1-1-4.pdf.

Wilburn:1926:LCR

[Wil26] S. D. Wilburn. Line current regulation in bridge polar duplex telegraph circuits. The Bell System Technical Journal, 5(4):625–635, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-625.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-625.pdf.

Wise:1929:ADR

[Wis29] W. Howard Wise. Asymptotic dipole radiation formulas. The Bell System Technical Journal, 8(4):662–671, October 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
4-662.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-4-662.pdf.

Wegel:1924:EFA

[WM24] R. L. Wegel and C. R. Moore.

An electrical frequency analyzer. The Bell System Technical Journal, 3(2):299–323, April 1924. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol03/bstj3-
2-299.pdf;   http://www.alcatel-
lucent.com/bstj/vol03-1924/articles/
bstj3-2-299.pdf.

Willis:1926:LCC

[WM26] F. C. Willis and L. E. Melhuish. Load carrying capacity of amplifiers. The Bell System Technical Journal, 5(4):573–592, October 1926. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol05/bstj5-
4-573.pdf;   http://www.alcatel-
lucent.com/bstj/vol05-1926/articles/
bstj5-4-573.pdf.

Williams:1929:PIM

[WM29] R. R. Williams and E. J. Murphy. The predominating influence of moisture and electrolytic material upon textiles as insulators. The Bell System Technical Journal, 8(2):225–242, April 1929. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol08/bstj8-
2-225.pdf;   http://www.alcatel-
lucent.com/bstj/vol08-1929/articles/
bstj8-2-225.pdf.

Wolfe:1925:CTH

[Wol25] W. V. Wolfe. Carrier telephony on high voltage power lines. The Bell System Technical Journal, 4(1):152–177, January 1925. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol04/bstj4-
1-152.pdf;   http://www.alcatel-
lucent.com/bstj/vol04-1925/articles/
bstj4-1-152.pdf.

Wright:1927:NYL

[WS27] S. B. Wright and H. C. Silent.

The New York–London telephone circuit. The Bell System Technical Journal, 6(4):736–749, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-736.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-736.pdf.

Wente:1928:HER

[WT28] E. C. Wente and A. L. Thuras. A high efficiency receiver for a horntype loud speaker of large power capacity. The Bell System Technical Journal, 7(1):140–153, January 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol07/bstj7-
1-140.pdf;   http://www.alcatel-
lucent.com/bstj/vol07-1928/articles/
bstj7-1-140.pdf.

Young:1927:PPT

[You27] R. L. Young. Power plants for telephone offices. The Bell System Technical Journal, 6(4):702–721, October 1927. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol06/bstj6-
4-702.pdf;   http://www.alcatel-
lucent.com/bstj/vol06-1927/articles/
bstj6-4-702.pdf.

Zobel:1923:TDU

[Zob23] Otto J. Zobel. Theory and design of uniform and composite electric wave-filters. The Bell System Technical Journal, 2(1):1–46, January 1923. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol02/bstj2-
                                        1-1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol02-1923/articles/
                                        bstj2-1-1.pdf.

Zobel:1924:TCE

[Zob24] Otto J. Zobel. Transmission characteristics of electric wavefilters. The Bell System Technical Journal, 3(4):567–620, October 1924. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol03/bstj3-
                                        4-567.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol03-1924/articles/
                                        bstj3-4-567.pdf.

Zobel:1928:DCE

[Zob28] Otto J. Zobel. Distortion correction in electrical circuits with constant resistance recurrent networks. The Bell System Technical Journal, 7(3):438–534, July 1928. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol07/bstj7-
                                        3-438.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol07-1928/articles/
                                        bstj7-3-438.pdf.